# It Couldn’t Help But Overhear: On the Limits of Modelling Meta-Communicative Grounding Acts with Supervised Learning

### v1

@@ 4 Variability in Human Grounding Acts
Backchannels, one form of positive evidence, have been demonstrated to involve individual variability, and even idiosyncrasy, which can be due to personality traits, gender or randomness (Huang and Gratch, 2012; Blomsma et al., 2024). The SotA for the task of backchannel prediction is not very high (.66 weighted F1) (Liermann et al., 2023).
Findings on the variability of human decisions to initiate a CR are still sparse. Stoyanchev et al. (2013) measured an absolute agreement of 39% among three annotators for scripted dialogues with missing ASR information. As another reference, Shaikh et al. (2023) reported a Cohen’s $\kappa$ of 48.45 for clarification in emotional support conversations, which, they claimed, may even be inflated. The task of deciding when to request clarification in collaborative instruction following is under active investigation, but models’ performance is still suboptimal (Shi et al., 2022; Li et al., 2023; Madureira and Schlangen, 2023; Mohanty et al., 2023). Recent works on the multimodal CoDraw dialogue game (Kim et al., 2019) argued that this may be due to the variability in human decisions and the limitations of using supervised learning (Testoni and Fernández, 2024; Madureira and Schlangen, 2024).

### v2

@@ 4 Variability in Human Grounding Acts
Backchannels, a positive evidence of grounding, were demonstrated to involve individual variability, and even idiosyncrasy, possibly due to personality, gender or randomness (Huang and Gratch, 2012; Blomsma et al., 2024). Although those works showed some regularity in their timing, the SotA for the backchannel prediction task is not very high (.66 weighted F1) (Liermann et al., 2023).
Findings on the variability of human decisions to initiate a CR, a negative sign of grounding, are still sparse. Stoyanchev et al. (2013) measured an absolute agreement of 39% among three annotators for scripted dialogues with missing ASR information. As another reference, Shaikh et al. (2023) reported a Cohen’s $\kappa$ of 48.45 for clarification in emotional support conversations, which, they claimed, may even be inflated. The task of deciding when to request clarification in collaborative instruction following is under active investigation, but models’ performance is still suboptimal (Shi et al., 2022; Li et al., 2023; Madureira and Schlangen, 2023; Mohanty et al., 2023). Recent works on the multimodal CoDraw dialogue game (Kim et al., 2019) argued that this may be due to the variability in human decisions and the limitations of using supervised learning (Testoni and Fernández, 2024; Madureira and Schlangen, 2024).